# Differentiable Logic Synthesis: Spectral Coefficient Selection via Sinkhorn-Constrained Composition

### v2

@@ 8.3 Hardware Deployment
8.4 Why Not SAT/ILP/Enumeration?
SAT solvers, integer linear programming (ILP), and brute-force enumeration are strong alternatives for discrete logic synthesis—and in fact, we use enumeration for $n=3$ precisely because it’s tractable. Our aim is not to claim superiority over these methods, but rather to demonstrate three complementary advantages:
Differentiable Integration.
SAT/ILP operate as black-box oracles: given a specification, they return a solution (or UNSAT). Our spectral approach enables gradient-based integration within larger differentiable systems. For instance, training an end-to-end neural architecture that includes learned Boolean logic blocks requires backpropagation through the logic layer—impossible with discrete solvers but natural in our Gumbel-Softmax formulation.
Warm-Start Capabilities.
The Phase 4 warm-start experiment (Section 7) demonstrates that exact spectral analysis (WHT) provides superior initialization for discrete refinement compared to random search. While SAT solvers excel at satisfiability, they lack the spectral structure that enables hierarchical decomposition and coefficient estimation for large $n$.
Compilation to Ternary PTFs.
Our synthesis targets a specific representation class: ternary polynomial threshold functions implementable as single-cycle combinational logic. SAT/ILP can verify logic equivalence but don’t inherently produce sparse spectral representations. The sparsity we achieve (31-39% across phases) reflects spectral concentration specific to the Walsh basis—structure that SAT encodings don’t directly exploit.
That said, hybrid approaches combining SAT-based verification with spectral coefficient search represent a promising direction for scaling beyond $n=4$, and we view these methods as complementary rather than competing paradigms.
@@ 9.1 Track 1: Exact FWHT (Moderate $n$)
9.2 Track 2: Coefficient Estimation (Large $n$)
For $n>28$, we estimate individual Fourier coefficients via Monte Carlo sampling from an oracle (black-box function). Given oracle access to $f:\{-1,+1\}^{n}\to\{-1,+1\}$:
@@ 9.2 Track 2: Coefficient Estimation (Large $n$)
By Hoeffding’s inequality, $m=O(\varepsilon^{-2}\log\delta^{-1})$ samples suffice for $|\hat{f}(S)-f(S)|\leq\varepsilon$ with probability $\geq 1-\delta$.
Search Strategy.
We restrict search to low-degree subsets (degree $\leq d$), motivated by the LMN theorem [4]: Fourier mass of depth-$d$ circuits concentrates on degree $\leq d$ terms. This reduces candidates from $2^{n}$ to $O(n^{d})$.
Honest Assessment.
This is a baseline Monte Carlo estimator, not a Goldreich-Levin style heavy coefficient finder. Query complexity is $O(n^{d}/\varepsilon^{2})$ for degree-$d$ search, compared to $\tilde{O}(k/\varepsilon^{2})$ for advanced algorithms recovering $k$ heavy coefficients [6, 5]. Implementation of bucket-splitting algorithms remains future work.
@@ 11 Conclusion
We introduced Hierarchical Spectral Composition, adapting the mHC framework [1] from LLM training stability to Boolean logic synthesis. By extending Sinkhorn-constrained routing with column-sign modulation, we achieve:
  • Scalability: Exact FWHT at 1.64B coeffs/sec ($n\leq 28$, 268M coefficients); hierarchical composition for 64-bit adders (6.5M verified samples) and 128-bit equality comparators
  • Hardware: 10,959 MOps/s on GPU, single-cycle FPGA inference
The key insight is that mHC’s Birkhoff polytope projection provides the stability foundation, while column-sign modulation adds the expressivity needed for discrete logic. For larger dimensions, we introduced two complementary approaches: spectral synthesis via MCMC refinement ($n=4$), and hierarchical composition from learned primitives (practical circuits). This combination scales from $n=2$ to large composed circuits with perfect accuracy, establishing a foundation for hardware-efficient neuro-symbolic logic synthesis.

### v3

@@ 9.1 Track 1: Exact FWHT (Moderate $n$)
9.2 Track 2: Oracle Learning with Spectral Filtering
For functions beyond enumerable truth tables ($n\approx 12$2020), we extract Boolean logic from black-box oracles via Fourier coefficient estimation. We implement and compare five methods to test whether spectral structure (Parseval normalization, symmetry constraints) or Sinkhorn projection improves coefficient recovery:
Method 1: Monte Carlo Baseline.
Direct sampling estimator:
@@ 9.2 Track 2: Oracle Learning with Spectral Filtering
Query complexity: $O(n^{d}/\varepsilon^{2})$ for degree-$d$ search (restrict to low-degree by LMN theorem [4]).
Method 2: Goldreich-Levin (GL).
Pairwise independent hashing [6] reduces variance via:
$\hat{f}(S)=\mathbb{E}_{a,x}[f(x)\cdot\chi_{S}(x)\cdot\chi_{\langle a,x\rangle\bmod 2}]$ (27)
where $a\sim\text{Uniform}(\{0,1\}^{n})$. Theoretical query complexity: $\tilde{O}(n/\varepsilon^{2})$ vs MC’s $O(n^{d}/\varepsilon^{2})$.
Method 3–5: Spectral Filtering.
Post-process GL estimates with:
  • GL+Parseval: Enforce $\sum\hat{f}^{2}=1$ (energy conservation)
  • GL+Parseval+Sym: Add symmetry constraints (for majority: equal degree-$k$ coeffs, odd-degree only)
  • GL+Sinkhorn: Birkhoff polytope projection on coefficient matrix (test if Sinkhorn helps denoising)
Experimental Design.
We test three function families at $n=16$ (degree $\leq 3$ search):
  • Parity (4-sparse): True support at degree-4 (outside search space)
  • Majority: Low-degree concentration (LMN theorem applies)
  • Comparator: Structured predicate (hierarchical composition)
Each method synthesizes a ternary PTF mask, evaluated on 100K test samples. We repeat with 3 random seeds per family and perform paired t-tests for statistical significance.
Results.
Table 15 shows accuracy means $\pm$ std across 3 seeds.
Table 15: Oracle Learning: Comparative Method Accuracy ($n=16$, degree $\leq 3$)
Method Parity Majority Comparator
MC $0.483\pm 0.011$ $0.724\pm 0.001$ $0.579\pm 0.001$
GL $0.479\pm 0.001$ $0.474\pm 0.000$ $0.464\pm 0.000$
GL+Parseval $0.479\pm 0.001$ $0.474\pm 0.000$ $0.464\pm 0.000$
GL+Parseval+Sym $0.479\pm 0.001$ $\mathbf{0.857\pm 0.001}$ $0.464\pm 0.000$
GL+Sinkhorn $0.487\pm 0.011$ $0.474\pm 0.000$ $0.464\pm 0.000$
Key Findings:
  1. 1.
    Symmetry >> Algorithms: GL+Parseval+Sym achieves 85.7% on majority (vs 72.4% MC, 47.4% GL), a +38.3% boost ($p<0.001$). Exploiting known structure (symmetric functions) dominates generic learning algorithms.
  2. 2.
    GL Underperforms MC: Contrary to theoretical expectations, GL performs worse than MC on majority (47.4% vs 72.4%, $p<0.001$) and comparator (46.4% vs 57.9%, $p<0.001$). This suggests threshold selection ($\tau=0.1$) or sample size (m=10m=10K per coeff) may be suboptimal for this regime.
  3. 3.
    Parseval Has No Effect: GL+Parseval = GL exactly ($\Delta=0$). Energy normalization alone does not improve recovery.
  4. 4.
    Sinkhorn Does Not Help Denoising: GL+Sinkhorn shows no improvement over GL or GL+Parseval ($p>0.05$ for all families). This validates Hypothesis B: Sinkhorn is useful for routing (composition, Phase 2), not coefficient denoising. The Birkhoff polytope serves different roles in different contexts.
Interpretation for Explainability.
These results reveal a fundamental insight for transparent-by-design AI: domain structure beats black-box learning. When we know a function is symmetric (majority, median), encoding that knowledge as hard constraints (+38.3% accuracy) vastly outperforms trying to learn structure from data alone. This aligns with the X-NeSy vision of combining symbolic knowledge with neural learning—the spectral coefficients are interpretable features (each $\hat{f}(S)$ has semantic meaning: "correlation with parity of variables SS"), and injecting symbolic priors (symmetry) makes the learned representation both more accurate and more explainable.
@@ 9.3 Track 3: Hierarchical Composition
9.4 Track 4: Routing Mechanism Comparison
[…]
We compare three routing mechanisms:
  • Sinkhorn: Column-stochastic projection via iterative normalization ($K=20$), as used in Phases 2–4.
  • Gumbel-STE: Gumbel-Softmax with straight-through estimator (hard one-hot forward, soft gradient backward), following Mind the Gap [8].
  • ReinMax: Sinkhorn projection with annealed entropy regularization in the loss ($\lambda_{\text{ent}}:0.1\to 0$), inspired by CardNN exploration strategies.
[…]
Table 17: Routing Mechanism Comparison on Mode-Switching Task ($n=2$, 4 frozen primitives). Sinkhorn achieves the best accuracy–interpretability tradeoff. Gumbel-STE exhibits the discretization gap discussed in [8]: hard routing commits early but cannot escape suboptimal assignments.
Router Accuracy Steps to 100% Sparsity Specialization Sign Coh. Stability
Sinkhorn 100.0% 100 0.46 0.06 0.96 0.99
ReinMax 100.0% 100 0.41 0.06 0.92 0.99
Gumbel-STE 88.5% 0.75 0.54 0.82 1.00
Key Findings.
  1. 1.
    Sinkhorn achieves the best accuracy–interpretability tradeoff. Both Sinkhorn and ReinMax reach 100% accuracy by step 100, with Sinkhorn attaining the highest sign coherence (0.96)—closest to the ternary $\{-1,0,+1\}$ values required for zero-loss quantization. This confirms the Phase 2 finding that Sinkhorn constraints enable exact discretization.
  2. 2.
    Gumbel-STE exhibits the discretization gap. Hard one-hot routing commits early (sparsity 0.75, specialization 0.54) but plateaus at 88.5% accuracy—the router locks onto a suboptimal primitive assignment and cannot recover. This empirically validates the “discretization gap” analyzed in Mind the Gap [8], where STE gradients provide biased updates that prevent convergence to the global optimum.
  3. 3.
    Entropy annealing rescues ReinMax. Without annealing ($\lambda_{\text{ent}}=0.1$ fixed), the entropy bonus dominates the loss (loss $\approx-0.36$), keeping routing completely diffuse (sparsity 0.21). Annealing $\lambda_{\text{ent}}:0.1\to 0$ alongside temperature allows early exploration followed by late sharpening, recovering performance comparable to Sinkhorn.
  4. 4.
    Soft routing outperforms hard routing for Boolean logic. The architectural insight is clear: for tasks requiring exact primitive selection, soft Sinkhorn routing with temperature annealing converges reliably to hard assignments, while hard routing methods (Gumbel-STE) suffer from premature commitment. This justifies the Sinkhorn design choice carried through Phases 2–4.
9.4.1 Spectral Analysis of Routing Dynamics
Beyond task accuracy, we analyze the spectral structure of the routing matrix $P\in\mathbb{R}^{3\times 4}$ during training via its singular value decomposition. Since $P$ maps 3 input features to 4 primitives, it has 3 singular values $\sigma_{1}\geq\sigma_{2}\geq\sigma_{3}\geq 0$. We track two derived quantities:
  • Spectral gap: $\Delta=1-\sigma_{2}/\sigma_{1}$, measuring how rank-1-like the routing is. $\Delta\to 1$ means all features route to one dominant primitive; $\Delta\to 0$ means routing distributes across multiple primitives.
  • Spectral entropy: $H_{\sigma}=-\sum_{i}p_{i}\log p_{i}$ where $p_{i}=\sigma_{i}^{2}/\sum_{j}\sigma_{j}^{2}$, measuring the effective dimensionality of the routing map. Higher $H_{\sigma}$ indicates more distributed spectral energy.
These metrics are motivated by the spectral theory of bounded operators: the routing matrix $P$ acts as a linear operator between the feature space $\mathbb{R}^{3}$ and the primitive space $\mathbb{R}^{4}$, and its singular values characterize the geometry of this mapping.
Table 18: Spectral trajectory of routing matrix $P$ during training. Sinkhorn maintains a smooth, monotonic trajectory (gradual spectral redistribution). Gumbel-STE exhibits chaotic oscillations in $\Delta$, indicating unstable routing dynamics. ReinMax tracks Sinkhorn’s trajectory as entropy regularization anneals to zero.
Sinkhorn Gumbel-STE ReinMax
Step $\Delta$ $H_{\sigma}$ $\Delta$ $H_{\sigma}$ $\Delta$ $H_{\sigma}$
0 0.95 0.03 0.94 0.03 0.95 0.02
100 0.37 0.60 0.76 0.21 0.51 0.49
500 0.26 0.65 0.58 0.42 0.59 0.42
2000 0.24 0.66 0.54 0.46 0.54 0.47
3300 0.20 0.67 0.68 0.31 0.45 0.55
5000 0.20 0.67 0.26 0.65 0.26 0.65
Observation 1: Routing Uncertainty Principle.
The Boolean uncertainty principle [3] states that for $f:\{-1,+1\}^{n}\to\mathbb{R}$:
$|\text{supp}(f)|\cdot|\text{supp}(\hat{f})|\geq 2^{n}$ (28)
A function cannot be simultaneously sparse in both the physical and spectral domains. We observe an analogous phenomenon for routing: the spectral entropy $H_{\sigma}$ of the converged routing matrix satisfies $H_{\sigma}\geq H^{*}$ where $H^{*}$ is determined by the task’s spectral complexity. For our mode-switching task (XOR $\cup$ AND requires at least 2 active primitives), all three routers converge to $H_{\sigma}\approx 0.65$$0.67$ at convergence, regardless of mechanism. This suggests a routing uncertainty bound: the minimum spectral entropy required to represent the target function class.
Observation 2: CFL-Like Stability Condition.
In numerical PDE, the Courant-Friedrichs-Lewy (CFL) condition requires that discretization not outpace the physical signal: $|v|\cdot\Delta t/\Delta x\leq C$. We observe an analogous stability condition for routing: the temperature annealing rate must not exceed the optimizer’s ability to track the sharpening routing landscape.
Sinkhorn’s spectral gap $\Delta(t)$ evolves monotonically ($0.95\to 0.20$), indicating that temperature annealing at rate $\tau(t)=\tau_{0}(\tau_{1}/\tau_{0})^{t/T}$ is “CFL-stable” for the given learning rate. Gumbel-STE violates this condition: hard one-hot routing imposes $\Delta\approx 1$ instantaneously (via $\operatorname{argmax}$), outpacing the gradient signal. The result is chaotic oscillations in $\Delta$ (ranging from 0.21 to 0.68 between steps 3000–3500) while accuracy remains trapped at 88.5%.
ReinMax’s entropy regularization acts as a CFL stabilizer: it constrains $\Delta$ from growing too fast early in training ($\Delta\leq 0.59$ for the first 1000 steps vs. Gumbel-STE’s $\Delta=0.76$ at step 100), then gradually releases this constraint as $\lambda_{\text{ent}}\to 0$, converging to the same stable endpoint as Sinkhorn.
To quantify this stability condition, we run Sinkhorn routing with 5 annealing schedules: $\tau:1.0\to 0.1$ over $T\in\{10{,}000,\;5{,}000,\;2{,}000,\;500\}$ steps plus an “instant” schedule ($\tau=0.01$ fixed). The annealing rate $r=\ln(\tau_{0}/\tau_{1})/T$ sweeps 4 orders of magnitude:
Table 19: CFL Stability Experiment: accuracy and spectral diagnostics vs. annealing rate $r$ for Sinkhorn routing (lr =0.01=0.01). A sharp transition between $r=0.0012$ and $r=0.0046$ separates stable convergence from failure.
Schedule $r$ Acc $\Delta_{\text{final}}$ $\text{Var}(\Delta)$ Status
Very Slow (T=10T{=}10K) $2.3\times 10^{-4}$ 100% 0.20 0.006 CFL-stable
Slow (T=5T{=}5K) $4.6\times 10^{-4}$ 100% 0.20 0.011 CFL-stable
Medium (T=2T{=}2K) $1.2\times 10^{-3}$ 100% 0.19 0.026 CFL-stable
Fast (T=500T{=}500) $4.6\times 10^{-3}$ 87.3% 0.23 0.076 CFL-violated
Instant ($\tau{=}0.01$) $\infty$ 100% 0.09 0.000 Alternate basin
The critical rate $r_{\text{crit}}\in(1.2\times 10^{-3},\;4.6\times 10^{-3})$ for lr =0.01=0.01 marks a sharp phase transition: below $r_{\text{crit}}$, the spectral gap evolves monotonically and accuracy reaches 100%; above it, the optimizer cannot track the sharpening loss landscape, producing $\text{Var}(\Delta)=0.076$ and 87.3% accuracy. The “instant” schedule ($\tau=0.01$ fixed) reaches a qualitatively different solution: $H_{\sigma}=1.06$ (vs. $\approx 0.67$ for annealed runs), $\sigma_{3}=0.82$ (vs. $\approx 0.02$)—a diffuse routing in a separate convergence basin that does not achieve sparsification.
Observation 3: Spectral Margin and Adaptability.
The smallest singular value $\sigma_{3}$ reveals a structural difference between routing mechanisms. At convergence:
  • Sinkhorn: $\sigma_{3}=0.024$, ReinMax: $\sigma_{3}=0.023$ (small but nonzero)
  • Gumbel-STE: $\sigma_{3}=9.2\times 10^{-9}$ (numerically zero)
Sinkhorn and ReinMax maintain a nonzero spectral margin—the routing is effectively rank-2 but retains a residual third singular direction. This margin serves as a “reserve capacity” for continued adaptation. Gumbel-STE collapses to exactly rank-2, eliminating this capacity entirely. The spectral margin thus provides a principled, information-theoretic metric for routing adaptability—complementing and grounding the empirical Gini-based sparsity measure used in prior sections.
Implications for Transparent-by-Design AI.
These spectral metrics offer a path toward provably interpretable routing:
  1. 1.
    The spectral gap $\Delta$ provides a single scalar measuring routing commitment, grounded in operator theory rather than ad hoc thresholds.
  2. 2.
    The spectral entropy $H_{\sigma}$ connects routing to information-theoretic bounds (uncertainty principle), enabling formal characterization of when routing is “sharp enough” for deployment.
  3. 3.
    The spectral margin $\sigma_{3}$ quantifies residual adaptability, distinguishing between “confidently committed” routing (low $\sigma_{3}$, stable $\Delta$) and “prematurely frozen” routing (zero $\sigma_{3}$, oscillating $\Delta$).
Together, these form an operator-theoretic interpretability framework for differentiable routing, moving beyond empirical metrics toward formally grounded transparency measures.
@@ 11 Conclusion
We introduced Hierarchical Spectral Composition, a transparent-by-design architecture for Boolean logic synthesis that makes learned representations intrinsically explainable through interpretable Fourier coefficients. By adapting the mHC framework [1] from LLM training stability to logic synthesis and extending Sinkhorn-constrained routing with column-sign modulation, we achieve:
  • n=2n=244: 100% accuracy on canonical Boolean operations via gradient descent ($n=2$), exhaustive search ($n=3$), and spectral synthesis with MCMC refinement ($n=4$), all converging to ternary masks $\{-1,0,+1\}$
  • Scalability: Exact FWHT at 1.64B coeffs/sec ($n\leq 28$, 268M coefficients); hierarchical composition for 64-bit adders and 128-bit equality comparators
  • Oracle Learning: Comparative evaluation of five coefficient estimation methods reveals that symbolic structure beats black-box learning—symmetry constraints boost majority accuracy +38% ($p<0.001$), while Sinkhorn projection aids routing but not coefficient denoising
  • Hardware: 10,959 MOps/s on GPU, single-cycle combinational logic inference
Explainability and Neurosymbolic Integration.
Our key contribution to explainable AI is demonstrating that spectral coefficients are interpretable features: each $\hat{f}(S)$ explicitly represents correlation with parity of variables $S$, making the learned logic transparent. The oracle learning experiments (Phase 5) establish a fundamental principle: injecting symbolic knowledge as hard spectral constraints (symmetry for majority, degree bounds for circuits) produces both higher accuracy and greater interpretability than generic learning algorithms. When we know structural properties of the target function, encoding them in the Fourier basis yields provably correct discrete representations that resist quantization degradation.
The Role of Birkhoff Projection.
Our comparative experiments resolve an open question: Sinkhorn’s Birkhoff polytope projection serves routing (Phase 2 composition) but not coefficient denoising (Phase 5 oracle learning). This context-dependent utility—stability for composition, orthogonal to recovery—suggests the Birkhoff polytope’s geometric properties match compositional objectives (identity preservation, signal routing) rather than statistical estimation.
Toward Hardware-Efficient Neuro-Symbolic Systems.
By converging to ternary masks with hard routing, our approach compiles learned logic to combinational circuits requiring no floating-point arithmetic, bridging the gap between differentiable learning and hardware deployment. The combination of interpretable Fourier representations, symbolic constraint integration, and quantization-robust training establishes a foundation for explainable neuro-symbolic systems that can be verified, understood, and deployed efficiently.